OpenStudy (anonymous):

a shopper purchased 4 tables and 2 chair for 200 and 2 tables and 7 chairs for 400 what is the cost of each table and each chair

OpenStudy (hoblos):

x=cost of table y=cost of chair 4x+2y =200 2x+7y=400 solve the system

OpenStudy (anonymous):

let the number of tables be x and number of chairs be y 4x+2y=200 ---(1) 2x+7y=400 ---(2) MUltiply (2) by 2, 4x+14y=800 (2) - (1) (subtract (1) from (2)) 4x+14y - 4x-2y= 800-200 12y= 600 y=50 Sub y=50 to (1) 4x+ 2(50) = 200 4x + 100=200 4x=100 x=25 Tables cost 25 dollars and chairs cost 50 dollars. Might i comment, how is the chair 2 times more expensive than the table.

OpenStudy (anonymous):

thank you